# Quantum mereology and subsystems from the spectrum

### v1

@@ V Discussion
  • At large scale it is easy to partition a system, systems are weakly interacting, observables supported on different subsystems almost commute, the ground state is localized and the physics of low energies and local observables can be well described in the diagonal basis,
@@
Finally, we suggest that the emergence of classicallity may be a tensor structure problem. A problem that consists in finding a tensor structure in which classical observables emerge and in which macroscopic subsystems behave classically. When a system is partitioned into weakly interacting sectors, slow evolving local quantities arise, e.g. the total energy of each sector. Furthermore, any operator $O$ that commutes with $H_{A}$ the Hamiltonian of a subsystem, almost commutes with the total Hamiltonian $H=H_{A}\otimes\mathbb{1}_{B}+\mathbb{1}_{A}\otimes H_{B}+H_{int}$, the error is just $[O,H_{int}]$. Finding slowly evolving local macroscopic quantities is crucial to understand the emergence of classicallity from quantum many body systems.

### v2

@@ 5 Discussion
  • At large scale it is easy to partition a system, systems are weakly interacting, observables supported on different subsystems almost commute (meaning that error $\epsilon$ in equation (3) is small), the ground state is localized and the physics of low energies and local observables can be well described in the diagonal basis,
@@
Finally, we suggest that the emergence of classicallity may be a tensor structure problem, as also mentioned in [7, 30, 2]. A problem that consists in finding a tensor structure in which classical observables emerge and in which macroscopic subsystems behave classically. When a system is partitioned into weakly interacting sectors, slow evolving local quantities arise, e.g. the total energy of each sector. Furthermore, any operator $O$ that commutes with $H_{A}$ the Hamiltonian of a subsystem, almost commutes with the total Hamiltonian $H=H_{A}\otimes\mathbb{1}_{B}+\mathbb{1}_{A}\otimes H_{B}+H_{int}$, the error is just $O,H_{int}]$. Finding slowly evolving local macroscopic quantities is crucial to understand the emergence of classicallity from quantum many body systems.
[…]
Code availability A Python package for quantum mereology is available at https://github.com/nicolasloizeau/quantum_mereology.
[…]
Acknowledgements N.L. was partially supported by a research grant (42085) from Villum Fonden. We thank Soren Schlassa, Lennart Fernandes, and Flaviano Morone for their discussions.